OpenStudy (anonymous):

x^2 - 2x =10 solve by completing the square use commas to separate answers as needed or use radicals as needed

OpenStudy (anonymous):

so you are going to take the coefficient on the x term and divide it by 2 than square it, than add it to both sides an simplify: \[x^2-2x=10\] so: \[\frac{-2}{2}=1 \rightarrow (1)^2=1\] \[x^2-2x+1=10+1\rightarrow x^2-2x+1=11\] The LHS can be factored, and so: \[(x-1)^2=11 \rightarrow x-1=\pm \sqrt{11} \rightarrow x=1\pm \sqrt{11}\]

OpenStudy (anonymous):

typo: \[\frac{-2}{2}=-1 \rightarrow (-1)^2=1\]